# High-dimensional quantum XYZ product codes for biased noise

### v2

@@ IV Numerical simulations
For instance, the bias rate of depolarizing and pure Pauli $Z$ noise model is $\eta=0.5$ and $\eta=\infty$, respectively. In this paper, we only consider noiseless stabilizer measurements.
@@
We should emphasize that, in this paper we do not explore an efficient and highly accurate decoder for 3D and 4D XYZ product codes, and FDBP we used here is an improved binary BP, which can be applied to all quantum stabilizer codes and have better decoding performance than tradition binary BP[21, 22], especially for some non-CSS codes[16]. However, as shown in Ref. [23], even though tradition binary BP-OSD[24] has been shown to have pretty good decoding performance for a wide range of QLDPC codes, but its performance can degrade for the codes with large code length and high degeneracy. FDBP also has similar problem and its decoding performance is also greatly degraded by short cycles, especially 4-cycles in the Tanner graph. In Appendix B, we compare the number of 4-cycles in the Tanner graphs of 3D Chamon code, 3D toric codes, 4D Chamon codes and 4D toric codes. Our computed result shows that the numbers of 4-cycles in the Tanner graph of 3D and 4D XYZ product codes are much more than those of 3D and 4D homological product codes. Thus, the error-correcting performance of 3D Chamon codes, 4D Chamon codes and 4D XYZ concatenated codes obtained by FDBP-OSD-0 in this paper is not the optimal. Nonetheless, from the perspective of code-capacity threshold, they still show good error-correcting performance for $Z$-biased noise model.
IV.1 3D Chamon code
@@
3D Chamon code proposed in Ref. [14] can be seen as 3D XYZ product of three repetition codes[9]. Bravyi et al. study some characteristics of this code in depth in Ref. [15]. However, its error-correcting performance has not been studied in depth so far.
@@
The dimension of 3D Chamon code is $4\gcd{\left(n_{1},n_{2},n_{3}\right)}$, where $n_{1}$, $n_{2}$ and $n_{3}$ are the code length of three repetition codes, respectively. Besides, we can also understand 3D Chamon code from the perspective of three-dimensional layout. As shown in Fig. 6, on a three-dimensional cubic Bravais lattice, the qubits of 3D Chamon code are placed on the face centers each of which neighbors two cubes and on the vertices each of which is shared in eight cubes. Each edge of cube represents a stabilizer that involves neighboring six qubits. These six qubits are divided into three classes, each of which contains two qubits along one of the $x$, $y$, $z$ three directions. For each class of qubits, the type of the operators that the stabilizer acts on them is Pauli $X$ or Pauli $Y$ or Pauli $Z$ according to the type of direction. We conjecture that 3D Chamon code may have the same error-correcting capability for Pauli $X$, $Z$ and $Y$ errors when $n_{1}=n_{2}=n_{3}$, since there exists spatial symmetry in this case.
@@
Figure 6: Three-dimensional layout of Chamon code. Qubits are placed on the face centers and the vertices of the cubes. Each edge of the cube represents a stabilizer, which involves neighboring six qubits. Some qubits are omitted for clearly.
@@
IV.2 4D Chamon code
We name the code which is 4D XYZ product of four repetition codes with code length $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ the 4D Chamon code and conjecture that it is 4D generalization of 3D Chamon code, just as the relation between 3D toric code and 4D toric code. There are two reasons. First, from the perspective of code dimension, the dimension of 4D Chamon code is $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, which is similar with $4\gcd{\left(n_{1},n_{2},n_{3}\right)}$, the dimension of 3D Chamon code. Second, from the perspective of code construction, the tensor structures of 3D and 4D XYZ product, as shown in Fig. 2 and 3, are the same as those of 3D and 4D homological product, respectively, while 3D and 4D toric codes are constructed by 3D and 4D homological product, respectively. We also conjecture that 3D and 4D Chamon codes may have similar topological structures, but in this paper we do not make further exploration for it.
In Fig. 8, we first consider the case of $n_{1}=n_{2}=n_{3}=n_{4}$. Fig. 8(a) and (b) show the single logical qubit error rate against the physical qubit error rate $p$ of 4D Chamon codes and 4D toric codes under depolarizing noise model, respectively. In this case, the 4D Chamon code encodes $8n_{1}^{2}$ logical qubits, while the 4D toric codes always encodes 6 logical qubits. In the legend, we use notation $L=n_{1},n_{2},n_{3},n_{4}$ to represent the corresponding code, which is constructed by four repetition codes with code length of $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$. One can see that the code-capacity threshold of 4D Chamon codes and 4D toric codes obtained by FDBP-OSD-0 is around $12.8\%$ and $15.9\%$, respectively, and the single logical qubit error rate of 4D Chamon codes is much higher than that of 4D toric codes. We attribute this to the fact that the number of 4-cycles of 4D Chamon codes is much more than that of 4D toric codes as shown in Appendix B, which degrades the decoding performance of FDBP-OSD-0.
@@
The dimension of 4D Chamon code is $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, when $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are coprime, the corresponding code always encode 8 logical qubits, which is close to 4D toric codes. Thus, it is more reasonable to compare the error-correcting performance between 4D Chamon codes and 4D toric codes in this case. First, we consider the pure Pauli $Z$ noise model. As shown in Fig. 9 (a) and (b), the code capacity of 4D Chamon codes and 4D toric codes obtained by FDBP-OSD-0 under Pauli $Z$ noise is around $18\%$ and $9.5\%$, respectively. This result means, form the perspective of code-capacity threshold, 4D Chamon codes have better error-correcting performance than that of 4D toric codes under pure Pauli $Z$ noise.
Figure 9: When $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are coprime, the error-correcting performance of (a) 4D Chamon codes and (b) 4D toric codes under pure Pauli $Z$ noise model.
@@
Figure 10: When $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are coprime, The error-correcting performance of 4D Chamon codes under $Z$-biased noise with bias rate (a) $\eta=1000$, (b) 100, (c) 10 and (d) 0.5.
IV.3 4D XYZ concatenated code
As described in Sect. III.2, 4D XYZ concatenated code is constructed from two concatenated code, which are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ ($n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd), and only encodes one logical qubit. As proved in Sect. III.3 the minimum weight of its logical $Z$ operator is $n_{3}n_{4}$, thus we conjecture that increasing block lengths $\left(n_{3},\ n_{4}\right)$ of the pair of repetition codes can construct the corresponding 4D XYZ product codes, which have better error-correcting performance for $Z$-biased noise than the codes constructed by 4D homological product.
@@
Figure 11: The error-correcting performance of (a) 4D XYZ concatenated codes and (b) 4D homological concatenated codes under pure $Z$ noise model.
Further, we explore the error-correcting performance of 4D XYZ concatenated codes under $Z$-biased noise with bias rate $\eta=1000$, 100, 10 and 0.5. Our simulation results show that the corresponding code-capacity threshold is around $37\%$, $32\%$, $22\%$ and $10\%$, respectively, which support that 4D XYZ product code can construct codes with good error-correcting performance for $Z$-biased noise.
Figure 12: The error-correcting performance of 4D XYZ concatenated codes under $Z$-biased noise with bias rate (a) $\eta=1000$, (b) 100, (c) 10 and (d) 0.5.
V Conclusion
@@
The results above mean that using the same two CSS codes, 4D XYZ product can construct non-CSS codes with better error-correcting performance for $Z$-biased noise than CSS codes constructed by 4D homological product, which is more meaningful for practice quantum computing system.
@@ Acknowledgements
Appendix A Proof of Lemma 2

### v3

@@ 4 Numerical simulations
For instance, the bias rate of depolarizing and pure Pauli $Z$ noise is $\eta=0.5$ and $\eta=\infty$, respectively. In this paper, we only consider noiseless stabilizer measurements.
@@
It is important to note that this paper does not focus on optimizing the decoding algorithm for 3D and 4D XYZ product codes. Instead, we use the FDBP-OSD-0 algorithm, which is an improved version of binary BP and is applicable to all quantum stabilizer codes with improved performance over traditional binary BP.[23, 24]. However, as shown in Ref. [25], even though tradition binary BP combined with OSD[26] has demonstrated strong decoding performance for a wide range of QLDPC codes, but its performance can degrade for the codes with large code length and high degeneracy. FDBP also has the same problem, and its performance is also greatly degraded by short cycles, especially 4-cycles in the Tanner graph. In Appendix C, we compare the number of 4-cycles in the Tanner graphs of 3D Chamon code, 3D toric codes, 4D Chamon codes and 4D toric codes. Our analysis reveals that the 4D XYZ product codes exhibit significantly more 4-cycles in their Tanner graphs compared to 3D and 4D homological product codes. Thus, the error-correcting performance of 3D Chamon codes, 4D Chamon codes and 4D XYZ product concatenated codes obtained by FDBP-OSD-0 in this paper is not the optimal. Nonetheless, from the perspective of code-capacity threshold, they still show good error-correcting performance against $Z$-biased noise.
4.1 3D Chamon code
@@
3D Chamon code proposed in Ref. [16] is an instance of 3D XYZ product of three repetition codes[5] and can be seen as a non-CSS variant of 3D toric code. Bravyi et al. study its characteristics in depth in Ref. [17]. However, its error-correcting performance under different noise models remains unexplored. Our work aims to address this gap by evaluating the code’s performance under various Pauli and depolarizing noise models.
@@
The code dimension of 3D Chamon code is $4\gcd{\left(n_{1},n_{2},n_{3}\right)}$, where $n_{1}$, $n_{2}$ and $n_{3}$ are the code length of three repetition codes, respectively. Besides, we can also understand 3D Chamon code from the perspective of three-dimensional layout. As shown in Fig. 6, on a three-dimensional cubic Bravais lattice, the qubits of 3D Chamon code are placed on the face centers each of which neighbors two cubes and on the vertices each of which is shared in eight cubes. Each edge of cube represents a stabilizer that involves neighboring six qubits. These six qubits are divided into three classes, each of which contains two qubits along one of the $x$, $y$, $z$ three directions. For each class of qubits, the type of the Pauli operators that the stabilizer acts on them is consistent with the type of direction. We conjecture that 3D Chamon code may have the same error-correcting capability for Pauli $X$, $Z$ and $Y$ errors when $n_{1}=n_{2}=n_{3}$, since there exists spatial symmetry in this case.
@@
Figure 6: Three-dimensional layout of 3D Chamon code. Qubits are placed on the face centers and the vertices of the cubes. Each edge of the cube represents a stabilizer, which involves neighboring six qubits. Some qubits are omitted for clearly.
@@
4.2 4D Chamon code
We refer the code which is 4D XYZ product of four repetition codes with code length $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ to 4D Chamon code and conjecture that it is 4D generalization of 3D Chamon code, just as the relation between 3D toric code and 4D toric code. There are two reasons. First, the code dimension of 4D Chamon code is $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, which is similar with $4\gcd{\left(n_{1},n_{2},n_{3}\right)}$, the code dimension of 3D Chamon code. Second, from the perspective of code construction, the tensor structures of 3D and 4D XYZ product, as shown in Fig. 2 and 3, are the same as those of 3D and 4D homological product, respectively, while 3D and 4D toric codes are constructed from 3D and 4D homological product, respectively.
In Fig. 8, we first consider the case of $n_{1}=n_{2}=n_{3}=n_{4}$. Fig. 8(a) and (b) show the single logical qubit error rate against the physical qubit error rate $p$ of 4D Chamon codes and 4D toric codes under depolarizing noise model, respectively. In this case, the 4D Chamon code encodes $8n_{1}^{2}$ logical qubits, while the 4D toric codes always encodes 6 logical qubits. In the legend, we use notation $L=n_{1},n_{2},n_{3},n_{4}$ to represent the corresponding code, which is constructed by four repetition codes with code length of $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$. One can see that the code-capacity threshold of 4D Chamon codes and 4D toric codes obtained by FDBP-OSD-0 is around $12.8\%$ and $15.9\%$, respectively, and the single logical qubit error rate of 4D Chamon codes is much higher than that of 4D toric codes. We attribute this to the fact that the number of 4-cycles of 4D Chamon codes is much more than that of 4D toric codes as shown in Appendix C, which degrades the decoding performance of FDBP-OSD-0.
@@
The code dimension of 4D Chamon code is $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, when $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are coprime, the corresponding code always encode 8 logical qubits, which is close to 4D toric codes. Thus, it is more reasonable to compare the error-correcting performance between 4D Chamon codes and 4D toric codes in this case. First, we consider the pure Pauli $Z$ noise model. As shown in Fig. 9 (a) and (b), the code capacity of 4D Chamon codes and 4D toric codes obtained by FDBP-OSD-0 under Pauli $Z$ noise is around $18\%$ and $9.5\%$, respectively. This result means, form the perspective of code-capacity threshold, 4D Chamon codes have better error-correcting performance than that of 4D toric codes against pure Pauli $Z$ noise.
Figure 9: When $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are coprime, the error-correcting performance of (a) 4D Chamon codes and (b) 4D toric codes against pure Pauli $Z$ noise.
@@
Figure 10: When $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are coprime, The error-correcting performance of 4D Chamon codes against $Z$-biased noise with bias rate (a) $\eta=1000$, (b) 100, (c) 10 and (d) 0.5.
4.3 4D XYZ product concatenated code
As described in Sect. 3.2, 4D XYZ product concatenated code is constructed from two concatenated code, which are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ ($n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd), and only encodes one logical qubit. According to Theorem 2 and Corollary 4, the minimum weight of its logical $Z$ operator is $n_{3}n_{4}$, thus we conjecture that increasing the block lengths $\left(n_{3},\ n_{4}\right)$ of the pair of repetition codes can construct the corresponding 4D XYZ product concatenated codes, which have better error-correcting performance against $Z$-biased noise than 4D homological product concatenated codes.
@@
Figure 11: The error-correcting performance of (a) 4D XYZ product concatenated codes and (b) 4D homological product concatenated codes against pure $Z$ noise.
Further, we explore the error-correcting performance of 4D XYZ product concatenated codes against $Z$-biased noise with bias rate $\eta=1000$, 100, 10 and 0.5. Our simulation results show that the corresponding code-capacity threshold is around $37\%$, $32\%$, $22\%$ and $10\%$, respectively, which support that 4D XYZ product code can construct codes with good error-correcting performance for $Z$-biased noise.
Figure 12: The error-correcting performance of 4D XYZ product concatenated codes against $Z$-biased noise with bias rate (a) $\eta=1000$, (b) 100, (c) 10 and (d) 0.5.
5 Conclusion
In this paper, exploiting FDBP-OSD-0, we first explore the error-correcting performance of 3D Chamon codes, which has not been studied in depth before. For 3D Chamon codes constructed by 3D XYZ product of three repetition codes with code length $n_{1}$, $n_{2}$, $n_{3}$, when $n_{1}=n_{2}=n_{3}$, simulation results show that the corresponding code-capacity threshold is around $14\%$ under depolarizing noise model and $13\%$, $14\%$ and $14\%$ under pure Pauli $X$, $Y$ and $Z$ noise model, respectively. Then we show that XYZ product can be generalized to four dimension and propose 4D XYZ product code construction, which constructs a class of non-CSS codes by using 4 classical codes or 2 CSS codes. Compared with 4D homological product, 4D XYZ product can construct non-CSS codes with higher code dimension or code distance.
@@ 5 Conclusion
These results indicate that, using the same two CSS codes, the 4D XYZ product can construct non-CSS codes with superior error-correcting performance against $Z$-biased noise compared to the CSS codes generated by the 4D homological product.
@@ Acknowledgements
Appendix A Proof of Theorem 2
Proof.
Our goal is to find out all independent logical operators, and our method is divided into three steps. First, finding out some operators that commutes with all stabilizer generators in Eq. (16). Second, excluding those operators which are stabilizers from the operators that we find in the first step, and the remaining operators are logical operators. Finally, we prove that the number of independent logical operator pairs that we find in the second step is equal to the code dimension of the corresponding 4D XYZ product code that we prove in Theorem 1.
If the form of $P_{1}$ is the same with that of $X_{G_{1}^{\prime}}$, we must have
[…]
$=\textbf{\emph{x}}\otimes H_{x_{2}}^{T}\textbf{\emph{i}}+\textbf{\emph{y}}\otimes H_{z_{2}}^{T}\textbf{\emph{j}}\in\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes Im\left(H_{x_{2}}^{T},H_{z_{2}}^{T}\right)$
$\begin{bmatrix}X_{G_{2}}\\ X_{G_{3}}\\ Z_{G_{2}}\\ Z_{G_{3}}\end{bmatrix}=\begin{bmatrix}X({\textbf{\emph{a}}_{1}^{\prime}}^{T})\ &I\ &I\ &Y({\textbf{\emph{b}}_{1}^{\prime}}^{T})\ &I\\ I\ &Y({\textbf{\emph{a}}_{2}^{\prime}}^{T})\ &I\ &I\ &X({\textbf{\emph{b}}_{2}^{\prime}}^{T})\\ Y({\textbf{\emph{c}}_{1}^{\prime}}^{T})\ &Z({\textbf{\emph{d}}_{1}^{\prime}}^{T})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z({\textbf{\emph{c}}_{2}^{\prime}}^{T})\ &Y({\textbf{\emph{d}}_{2}^{\prime}}^{T})\end{bmatrix}$ (39)
Let ${\pi\left({\textbf{\emph{a}}_{1}^{\prime\prime}}^{T},\ {\textbf{\emph{a}}_{2}^{\prime\prime}}^{T},{\textbf{\emph{b}}_{1}^{\prime\prime}}^{T},{\textbf{\emph{b}}_{2}^{\prime\prime}}^{T}\right)}^{T}\in\ker{\left(\left[H_{z1}^{T},H_{x1}^{T}\right]\right)}\otimes Im\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ and $\begin{bmatrix}X_{G_{2}}^{\prime}\\ X_{G_{3}}^{\prime}\end{bmatrix}=\begin{bmatrix}X({\textbf{\emph{a}}_{1}^{\prime\prime}}^{T})\ &I\ &I\ &Y({\textbf{\emph{b}}_{1}^{\prime\prime}}^{T})\ &I\\ I\ &Y({\textbf{\emph{a}}_{2}^{\prime\prime}}^{T})\ &I\ &I\ &X({\textbf{\emph{b}}_{2}^{\prime\prime}}^{T})\end{bmatrix}$. Next, we prove that $X_{G_{2}^{\prime}}X_{G_{3}^{\prime}}=\left[X\left({\textbf{\emph{a}}_{1}^{\prime\prime}}^{T}\right)\ Y\left({\textbf{\emph{a}}_{2}^{\prime\prime}}^{T}\right)\ I\ Y\left({\textbf{\emph{b}}_{1}^{\prime\prime}}^{T}\right)\ X\left({\textbf{\emph{b}}_{2}^{\prime\prime}}^{T}\right)\right]$ is a stabilizer, which can only be generated by the product of some stabilizer generators in $S$ and $V$, thus should be excluded. Considering an operator $P_{2}$ which is the product of some stabilizer generators in $S$ and $V$, it must have the following form
$P_{2}$ $=\left[\textbf{\emph{s}}^{T},\textbf{\emph{v}}^{T}\right]\begin{bmatrix}X^{\left(I_{m_{1}}\otimes H_{x_{2}}^{T}\right)}&Y^{\left(I_{m_{1}}\otimes H_{z_{2}}^{T}\right)}&Z^{\left(H_{z_{1}}\otimes I_{n_{B}}\right)}&I&I\\ I&I&Z^{\left(H_{x1}\otimes I_{n_{B}}\right)}&Y^{\left(I_{m_{2}}\otimes H_{x_{2}}^{T}\right)}&X^{\left(I_{m_{2}}\otimes H_{z_{2}}^{T}\right)}\end{bmatrix}$ (40)
If the form of $P_{2}$ is the same with $X_{G_{2}^{\prime}}X_{G_{3}^{\prime}}$, we must have
Thus, we have $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}\in\ker\left(\left[H_{z1}^{T}\otimes I_{n_{B}},H_{x1}^{T}\otimes I_{n_{B}}\right]\right)=\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\mathcal{C}_{n_{B}}$. Let $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}=\begin{bmatrix}\textbf{\emph{x}}\\ \textbf{\emph{y}}\end{bmatrix}\otimes\textbf{\emph{i}}$, where $\begin{bmatrix}\textbf{\emph{x}}\\ \textbf{\emph{y}}\end{bmatrix}=\left[x_{1},\cdots,x_{m_{1}},y_{1},\cdots,y_{m_{2}}\right]^{T}\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}$ and $\textbf{\emph{i}}\in\mathcal{C}_{n_{B}}$, we have
$\begin{bmatrix}I_{m_{1}}\otimes H_{x_{2}}&\textbf{0}\\ I_{m_{1}}\otimes H_{z_{2}}&\textbf{0}\\ \textbf{0}&I_{m_{2}}\otimes H_{x_{2}}\\ \textbf{0}&I_{m_{2}}\otimes H_{z_{2}}\end{bmatrix}\begin{bmatrix}\textbf{\emph{x}}\\ \textbf{\emph{y}}\end{bmatrix}\otimes\textbf{\emph{i}}=\begin{bmatrix}\textbf{\emph{x}}\otimes H_{x_{2}}\textbf{\emph{i}}\\ \textbf{\emph{x}}\otimes H_{z_{2}}\textbf{\emph{i}}\\ \textbf{\emph{y}}\otimes H_{x_{2}}\textbf{\emph{i}}\\ \textbf{\emph{y}}\otimes H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}=\begin{bmatrix}x_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{x_{2}}\textbf{\emph{i}}\\ x_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{z_{2}}\textbf{\emph{i}}\\ y_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{x_{2}}\textbf{\emph{i}}\\ y_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}$ (42)
It can be seen that
$\pi\left(\begin{bmatrix}x_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{x_{2}}\textbf{\emph{i}}\\ x_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{z_{2}}\textbf{\emph{i}}\\ y_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{x_{2}}\textbf{\emph{i}}\\ y_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}\right)=\begin{bmatrix}x_{1}H_{x_{2}}\textbf{\emph{i}}\\ x_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{x_{2}}\textbf{\emph{i}}\\ x_{m_{1}}H_{z_{2}}\textbf{\emph{i}}\\ y_{1}H_{x_{2}}\textbf{\emph{i}}\\ y_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{x_{2}}\textbf{\emph{i}}\\ y_{m_{2}}H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\otimes Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}$ (43)
which means $X_{G_{2}^{\prime}}X_{G_{3}^{\prime}}$ is a stabilizer. Thus, for vectors ${\pi\left(\textbf{\emph{a}}_{1}^{T},\textbf{\emph{a}}_{2}^{T},\textbf{\emph{b}}_{1}^{T},\textbf{\emph{b}}_{2}^{T}\right)}^{T}\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$, the corresponding operators $\begin{bmatrix}X_{L_{2}}\\ X_{L_{3}}\end{bmatrix}=\begin{bmatrix}X({\textbf{\emph{a}}_{1}}^{T})\ &I\ &I\ &Y({\textbf{\emph{b}}_{1}}^{T})\ &I\\ I\ &Y({\textbf{\emph{a}}_{2}}^{T})\ &I\ &I\ &X({\textbf{\emph{b}}_{2}}^{T})\end{bmatrix}$ are logical operators.
Similarly, for vector $\left({\textbf{\emph{c}}_{1}^{\prime\prime}}^{T},{\textbf{\emph{d}}_{1}^{\prime\prime}}^{T},{\textbf{\emph{c}}_{2}^{\prime\prime}}^{T},{\textbf{\emph{d}}_{2}^{\prime\prime}}^{T}\right)^{T}\in Im\left(\begin{bmatrix}H_{z_{1}}\\ H_{x_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)$ and $\begin{bmatrix}Z_{G_{2}}^{\prime}\\ Z_{G_{3}}^{\prime}\end{bmatrix}=\begin{bmatrix}Y({\textbf{\emph{c}}_{1}^{\prime\prime}}^{T})\ &Z({\textbf{\emph{d}}_{1}^{\prime\prime}}^{T})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z({\textbf{\emph{c}}_{2}^{\prime\prime}}^{T})\ &Y({\textbf{\emph{d}}_{2}^{\prime\prime}}^{T})\end{bmatrix}$, the operator $Z_{G_{2}^{\prime}}Z_{G_{3}^{\prime}}=\left[Y\left({\textbf{\emph{c}}_{1}^{\prime\prime}}^{T}\right)\ \ Z\left({\textbf{\emph{d}}_{1}^{\prime\prime}}^{T}\right)\ \ I\ \ Z\left({\textbf{\emph{c}}_{2}^{\prime\prime}}^{T}\right)\ \ Y\left({\textbf{\emph{d}}_{2}^{\prime\prime}}^{T}\right)\right]$ is a stabilizer which can only be generated by the product of some stabilizer generators in $T$ and $U$ and thus should be excluded.
To sum up, for vectors ${\pi\left(\textbf{\emph{a}}_{1}^{T},\ \textbf{\emph{a}}_{2}^{T},\textbf{\emph{b}}_{1}^{T},\textbf{\emph{b}}_{2}^{T}\right)}^{T}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ and ${\pi\left(\textbf{\emph{c}}_{1}^{T},\ \textbf{\emph{d}}_{1}^{T},\textbf{\emph{c}}_{2}^{T},\textbf{\emph{d}}_{2}^{T}\right)}^{T}\in\left(\mathcal{C}_{m_{1}+m_{2}}\backslash Im\left(\begin{bmatrix}H_{z_{1}}\\ H_{x_{1}}\end{bmatrix}\right)\right)\otimes\ker\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)$, the corresponding $\begin{bmatrix}X_{L_{2}}\\ X_{L_{3}}\end{bmatrix}=\begin{bmatrix}X(\textbf{\emph{a}}_{1})\ &I\ &I\ &Y(\textbf{\emph{b}}_{1})\ &I\\ I\ &Y(\textbf{\emph{a}}_{2})\ &I\ &I\ &X(\textbf{\emph{b}}_{2})\end{bmatrix}$ and $\begin{bmatrix}Z_{L_{2}}\\ Z_{L_{3}}\end{bmatrix}=\begin{bmatrix}Y(\textbf{\emph{c}}_{1})\ &Z(\textbf{\emph{d}}_{1})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z(\textbf{\emph{c}}_{2})\ &Y(\textbf{\emph{d}}_{2})\end{bmatrix}$ must be logical operators.
Notice that the dimension of vector space $\mathcal{C}_{m_{3}+m_{4}}\backslash Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}$ is $m_{3}+m_{4}-\dim\left(Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)=m_{3}+m_{4}-dim\left(row\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)=\dim\left(\ker\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$. Similarly, the dimension of vector space $\mathcal{C}_{m_{1}+m_{1}}\backslash Im\begin{bmatrix}H_{z_{1}}\\ H_{x_{1}}\end{bmatrix}$ is $m_{1}+m_{2}-\dim\left(Im\begin{bmatrix}H_{z_{1}}\\ H_{x_{1}}\end{bmatrix}\right)=m_{1}+m_{2}-dim\left(row\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\right)=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\right)$. Thus, the total number of independent logical operators $X_{L_{2}}$ and $X_{L_{3}}$ and that of $Z_{L_{2}}$ and $Z_{L_{3}}$ are both $\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$. Moreover, for any vector $\textbf{\emph{r}}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$, we can find a vector $\textbf{\emph{w}}\in\left(\mathcal{C}_{m_{1}+m_{2}}\backslash Im\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)$, such that $\textbf{\emph{r}}\cdot\textbf{\emph{w}}^{T}=1$, which means the corresponding $X_{L}$ and $Z_{L}$ anti-commute. Thus, there are $\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker{\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)}\right)$ pairs of the second type logical operators.
The total number of the first and the second types of logical operators is $\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\times\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)+\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker{\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)}\right)$. Now, we prove that this number is equal to the code dimension of the corresponding 4D XYZ product code as we prove in Theorem 1.
$\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\times\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)+\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker{\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)}\right)$ (44)
$=a\left[b-\left(n_{B}-m_{3}-m_{4}\right)\right]+b\left[a-\left(n_{A}-m_{1}-m_{2}\right)\right]+\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)$
which is equal to the code dimension as we prove in Theorem 1. Thus, the proof is completed. ∎
Appendix B Proof of Lemma 2